OpenStudy (anonymous):

whats the prime factorization for 200

OpenStudy (anonymous):

\[ 200 = 2^3 \times 5^2 \]

hero (hero):

You're probably wondering how he figured this out

hero (hero):

Basically, what you do is, start with 2. Keep dividing by two until you can't anymore. Then move to the next prime number, 3. If it's not divisible by three, move on to the next prime number, 5. Keep dividing by 5 until you no longer can. Once you get to 1, you're done.

myininaya (myininaya):

how many quarters are in 2 dollars 8 quarters right 8=2^3 so you can say 2^3 quarters

myininaya (myininaya):

so 200=2^3*25=2^3*5^2